OpenStudy (elleblythe):

Determine all values of k such that the graph of 2kx^2+2kx+k+1=0 intersects the x-axis exactly once. Please show complete solution and answer

sam (.sam.):

b2 - 4ac when it is zero you get just ONE real solution so b^2-4ac=0 What is a, b and c?

OpenStudy (elleblythe):

@.Sam. i tried solving for a, b and c and ended up with -4K^2-8k=0 what do i do next?

OpenStudy (elleblythe):

@.Sam. sorry i meant i tried solving for k, not a, b and c

sam (.sam.):

Yeah, just solve for k, add both sides -4k^2, 4k^2=-8k k=-2

OpenStudy (elleblythe):

@.Sam. isn't it supposed to be completing the square? how did you get 2 if there were k's on both sides of the equation?

sam (.sam.):

You divide both sides by k